Calculate the temperature (in ºC) to which silver ore, Ag2O, must be heated to spontaneously decompose the ore to oxygen gas and solid silver metal at standard state. (Δf of Ag2O(s) = –31.05 kJ/mol;   values: Ag(s), 42.55 J/mol•K; O2(g), 205.1 J/mol•K; Ag2O(s), 121.3 J/mol•K)

234 ºC
843 ºC
374 ºC
467 ºC
195 ºC
Answer #1

Ag2O(s) ------------> 2Ag(s) + 1/2O2(g)

ΔHº     = Δf products - Δf reactants

          = 1/2*0 + 2*0 - (-31.05)    = 31.05KJ/mole

ΔSº     = f products - f reactants

             = 1/2*205.1 + 2*42.55 - (121.3)

              = 102.55 +2*42.55 - (121.3)   = 66.35J/mole   = 0.06635KJ/mole-K

ΔGº    = ΔHº -TΔ

ΔGº <0 for spontaneous reaction

ΔGº <0

ΔHº -TΔSº <0

ΔHº <TΔ

ΔHº /ΔSº <T

31.05/0.06635<T

468K<T

1950C<T

1950C >>>>>answer